OpenStudy (anonymous):

|x+5| - 2|x| = 1 This shoul be easy but I'm stuck and need some help

OpenStudy (anonymous):

Hmm, yeah, the ones with two absolute value expressions can be tricky. You have to consider all the possibilities of what might be in the ||'s

OpenStudy (anonymous):

yeah it is not that easy. you have to check cases

OpenStudy (anonymous):

e.g. (x+5) and (x) could both be positive, both negative, one pos. the other neg. etc.

OpenStudy (anonymous):

yea, I'm aware of this

OpenStudy (anonymous):

maybe it would be best to start with \[y=|x+5|\] and \[y=2|x|+1\] and see where they intersect

OpenStudy (anonymous):

Then write the four different equations and solve each independently, then go back to the original statement to check the solutions.

OpenStudy (anonymous):

^^ good tip.

OpenStudy (anonymous):

first one is \(y=x+5,x\geq -5, y=-x-5, y<5\) second one is \(2x+1\) or \(-2x+1\) depending on whether \(x\geq 0\) or \(x< 0\) on one of these intervals there will be no solution i believe